OpenStudy (anonymous):

Eigenvalues and Eigenvectors... Find 1-dimensional subspaces of R2 invariant under the linear transformations given by the following matrix: [1 2;-3 -6]

OpenStudy (anonymous):

so you want the vectors v of the form: \[A\vec{v} = \lambda \vec{v}\] for the above matrix (which we call A) and some constant lambda. Do you know how to solve problems like this?

OpenStudy (anonymous):

Or you could just do it by looking at it, as 2x2 matrices aren't too hard to see through. Up to you.

OpenStudy (anonymous):

only just scratched he surface of eigenvalues/vectors but from my notes you rearrange \[Av =\lambda v\] to \[(A-\lambda I)v=0\] then find \[\det(A-\lambda I)=0\] and I get \[\lambda = 0, -5\] so what do I do now

OpenStudy (anonymous):

The determinant is \[(1-\lambda)( -6 - \lambda) +6 = -6 -\lambda + 6\lambda + \lambda^2 + 6 = \lambda^2 + 5\lambda \] so you're correct. The matrix that came from is \[\left[\begin{matrix}1-\lambda & 2\\-3 &-6- \lambda\end{matrix}\right]\] Plug in lambda = 0, and solve \[\left[\begin{matrix}1 & 2\\-3 &-6\end{matrix}\right]\left(\begin{matrix} a \\b\end{matrix}\right)= 0\] You get two equations. First, \[a + 2b = 0\] second, \[-3a - 6b = 0\] these happen to give the same results, namely a = -2b, and so therefore your eigenvector is \[\left(\begin{matrix}-2 \\ 1\end{matrix}\right)\] and its multiples.

OpenStudy (anonymous):

Ohh ok thanks, I haven't quite covered the eigenvectors yet but this makes complete sense! there's also another eigenvector [1 -3] which by the look of it you plug in the 5 as lambda in your 2nd step and solve. Thanks again!

OpenStudy (anonymous):

Exactly. Therefore, the subspaces defined by those two vectors and their multiples are invariant under the transformation.

OpenStudy (anonymous):

Just out of curiosity, what was you other method you described earlier, about reading off the 2x2 matrix?

OpenStudy (anonymous):

I just meant looking at the thing and trying to see without doing any computation. It's not so bad sometimes with 2x2 matrices but it's nearly impossible any other time.